A contractor is trying to estimate the distance to be expected of drawdown of 4.81 m from pumping well under the following conditions:

a) Pumping rate 0.028 m3/s

b) Pumping time = 1066 days

c) Drawdown in observation well = 9.52 m

d) Observation well is located 10.0 m from the pumping well

e) Aquifer material = medium sand

f) Aquifer thickness = 14.05 m

Assume that the well is fully penetrating in an unconfined aquifer.


1
Expert's answer
2022-03-24T01:40:46-0400

K = 1.4 "\times" 10-4m "\times" s-1


h1 = 14.05 - 9.52 = 4.53 m (h1)2 = 20.52 


h2 = 14.05 - 4.81 = 9.24 m (h2)2 = 85.38


"\frac{r2}{10.0}" = 10.0"\times" 2.7699


r2 = 27.70m